What rate of change of current in a solenoid having self-inductance 19.4 mH
produces a self-induced emf of 100 mV in it?
1
Expert's answer
2020-03-09T10:54:43-0400

"E_s=-L\frac{dI}{dt}"


"\lvert\frac{dI}{dt}\rvert=\frac{E_s}{L}=\frac{0.1}{0.0194}=5.15A/s" Answer